OpenStudy (anonymous):

prove using epsilon and delta that the limit of [x/2] as x approaches 3 =1

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} \lfloor x/2 \rfloor=1\]

OpenStudy (anonymous):

\[ \left\lfloor \frac{x}{2}\right\rfloor =1, \text { for } 1\le x \le 4 \] near 3 it is equal to 1 Let \( \epsilon >0\), take \(\delta =\frac 1 2\), then if \[| x -3| < \delta \implies \left | \left\lfloor \frac{x}{2}\right\rfloor -1 \right |=|1-1|=0 <\epsilon \]

OpenStudy (anonymous):

"let delta be a 1/2 " that was randomly picked number yes? Second question : how did you get to x-3?

OpenStudy (anonymous):

I think it's for \[1\le x <4\] and not including 4. Then I understand.

OpenStudy (anonymous):

yes, you are right.